A sample of HBr is introduced at one end of a 1-meter tube. A sample of gas with a molecular weight of 80.9 is introduced at the other. Use Graham's law to calculate where (how many meters from the HBr end) the two gases meet.
Answer #1

distnace travelled by HBr = x meters

unknown gas distnace travelled = 1 - x meters

rate of diffusion inversely proportional to square root of molar mass

R1 / R2 = sqrt (M2 / M1)

R1 / R2 = sqrt (80.9 / 80.9) = 1

R1 = R2

rate of diffusion is equal to distance travelled by time

r / t = r/ t

x = x -1

2 x = 1

x = 1/2

distnace travelled by HBr = 1/2 m

unknown gas distnace travelled = 1 - x = 1/2 m

they will meet at 1/5 meter of tube (half distance)